OpenStudy (anonymous):

the probability of certain hockey player making a goal after hitting a slap shot is 1/5. how many successful slap shots would you expect her to make after 120 attempts 5 20 24 60

OpenStudy (warriorz13):

1/5=.2

OpenStudy (anonymous):

Ok 1/5 to ?/120 what is the blank there

OpenStudy (anonymous):

idk

OpenStudy (warriorz13):

SO... 1/5=.2 .2*120=24?

OpenStudy (anonymous):

\[Total=Attempts \times Probability\]\[T=120 \times \frac{1}{5}\]\[T=\frac{120}{5}\]\[T=24\] We would expect her to score 24 goals.

OpenStudy (anonymous):

How is 1/5 2?

OpenStudy (texaschic101):

he makes 1 slapshot out of 5 shots.....120/5 = your answer

OpenStudy (anonymous):

Exactly texas

OpenStudy (warriorz13):

NONONONONONONO @IsaiahCC100 .2

OpenStudy (anonymous):

so 5 is my answer

OpenStudy (warriorz13):

So the answer is 24

OpenStudy (warriorz13):

DO .2*120

OpenStudy (anonymous):

\[120 \times \frac{1}{5}=120 \times 0.20=24\]

OpenStudy (texaschic101):

he makes 20%...so .20(120) = 24 you can do it either way

OpenStudy (warriorz13):

Which is 24

OpenStudy (anonymous):

thanks

OpenStudy (warriorz13):

Yes thank you!!!

OpenStudy (warriorz13):

NP See I can do it right LMFAO finally

OpenStudy (texaschic101):

yes you can :)